7. The probability that a pregnancy results in twins is about 0.03. The probability that a pregnancy results in identical twins is about 0.002. Given that a pregnancy results in twins, what is the probability they are identical twins? 

a) Write this question using the proper symbols.   

 

b) Find the answer. Round to the nearest ten thousandths place.  

 

 

 

 

8. If the events A and B are independent, P(A) = 0.62 and P(B) = 0.23, find P(A and B).


1
Expert's answer
2021-05-10T14:20:45-0400

7.

P(twins) = 0.03

P(identical twins) = 0.002

a) P(identical twins/twins)

b) P(identical twins/twins) "= \frac{0.002}{0.03} = 0.0667"

8.

"P(A) = 0.62 \\

P(B) = 0.23 \\

P(A∩B) = P(A) \times P(B) \\

= 0.62 \times 0.23 = 0.142 \\

P(A \cup B) = P(A ) + P(B) - P(A \cap B) \\

=( 0.62 + 0.23 ) - 0.142 = 0.708 \\

P(A \cup B) = 0.708"